what is the heigth of an equalateral triangle that is 30 ft wide by 30 ft on each side and formula to solve this problem?
Found 2 solutions by edjones, solver91311:
Answer by edjones(7311)

The height bisects the base of an equilateral triangle.
b=15 c=30
a^2+b^2=c^2
a^2+225=900
a^2=675
a=sqrt(3^3*5^2) prime factors of 675.
=15sqrt(3)'
.
Ed

Answer by solver91311(12126)



The altitude of an equilateral triangle is also a perpendicular bisector of the base. Therefore, the height is the measure of one leg of a right triangle where the hypotenuse is the measure of one side of the equilateral triangle and the other leg measures one-half of the side of the equilateral triangle.

So, applying Pythagoras, given the measure of the side, , the height, is:



Hence, for a 30 foot equilateral triangle, the height is:




John